129 ml of water are produced from the reaction of 123g of Iron (III) hydroxide and phosphorus acid. If the Iron (III) hydroxide was the limiting reagent, what would be the minimum concentration that 560 ml of acid would have?
Answer #1

The reaction that occurs is:

Fe (OH) 3 + H3PO3 = FePO3 + 3 H2O

The moles of H3PO3 used to react with Fe (OH) 3 are calculated:

n H3PO3 = 123 g Fe (OH) 3 * (1 mol Fe (OH) 3/107 g) * (1 mol H3PO3 / 1 mol Fe (OH) 3) = 1.15 mol

The minimum concentration is calculated:

[H3PO3] = n / V = ​​1.15 mol / 0.56 L = 2.05 M

If you liked the answer, please rate it in a positive way, you would help me a lot, thank you.